# Cosmic Ray Diffusion in the Turbulent Interstellar Medium: Effects of Mirror Diffusion and Pitch Angle Scattering

### v1

@@ 3.1 Particle trajectory and mirroring identification
In turbulence particle trajectories are distorted by scattering. To distinguish such events from mirror diffusion, we introduce the following criterium. Let us recall that $\mu_{b}$ is the peak value of $\mu$ right before a bounce. We can thus evaluate the respective adiabatic invariant of $\mu_{b}$, $\zeta$ using Equation 3. This is shown in the bottom panel of Figures 3 and 4. For a particle with $n$ occurrences of $\mu=0$, and each peak occurring at times $t_{1}<t_{2}<...<t_{i}<...<t_{n}$, we compare $\zeta(t_{i})/\zeta_{0}$ with $\zeta(t_{i-1})/\zeta_{0}$ and $\zeta(t_{i+1})/\zeta_{0}$ (to simplify notation, consider $\zeta_{i}=\zeta(t_{i})/\zeta_{0}$). With this, for a given bounce at time $t_{i}$, $\mu_{b}(t_{i})$ is discarded if $|\zeta_{i}-\zeta_{i-1}|>0.2$ and $|\zeta_{i}-\zeta_{i+1}|>0.2$. In other words, if the variation of $\zeta_{i}$ is bigger than $0.2$ compared to its neighbors, the bounce is discarded [^2] The value of 0.2 was arbitrarily chosen to ensure that we had a maximum of $\sim 20\%$ variation in the adiabatic invariant for each identified $\mu_{b}$.. For the first identified point, $\zeta_{1}$, we compare with the values of the next two identified points whilst for the last one, $\zeta_{n}$, we compare with the two before. If there are less than 3 bounce events, we simply accept the values of identified $\mu_{b}$.
@@ 3.2 Pitch angle distribution of $\mu_{b}$
Figure 5: Comparison between the expected values of $\mu_{c,max}$ (Equation 6) and the identified bounce values from the simulations $\mu_{b}$. The orange line indicates where $\mu_{c,max}=\mu_{b}$. The color map shows the density of points in the diagrams. Top panels show the results for models with the complete magnetic field distribution, while the bottom panels show the results for models with the smoothed distribution of magnetic fields. On the left hand side we show models with $\mu_{0}=0.2$, while on the right hand side we show models with $\mu_{0}=0.8$.
Figure 6: Probability density distributions of $\mu_{b}$ (see text for further details). Colors indicate different Larmor radii of the particles. The top panels show the PDFs for models with the complete background magnetic field, whilst the bottom panels show PDFs for models with the smoothed magnetic field as background. On the left hand side we show models with $\mu_{0}=0.2$, while on the right hand side we show models with $\mu_{0}=0.8$.
@@
Figure 5 compares values of $\mu_{b}$ with the respective $\mu_{c,max}$ for every bounce identified in the models with $R_{L}=0.03L_{0}$. The top panels show the results for models with complete distribution of magnetic fields, while the bottom panels show the same for models that consider the smoothed distribution of magnetic fields. When $\mu_{0}$ is small, the number of bounces is generally higher than when $\mu_{0}$ is closer to 1. Comparing the upper and bottom panels, it is clear that the lack of scattering plays an important role in slowing down the diffusion parallel to the magnetic field, as evidenced by the models with a smoothed background.
@@
In the smoothed models in Figure 5 (panels (c) and (d)) we see that the identified $\mu_{b}$ events are mostly smaller than the corresponding predicted maximum values $\mu_{c,max}$. This means that the particles are bouncing from larger amplitude mirrors in the MHD cascade. As shown in Figure 4, the peaks maintain similar values while the particle bounces through successive mirrors. Also, $\zeta$ stays almost constant during this process, which indicates that $\mu_{b}$ is a good proxy for the bouncing process. The probability of a bounce occurring for a certain value of $\mu$ is not very sensitive to the initial conditions, unless scattering is suppressed. As evidenced by panels (c) and (d) of Figure 5, where we have a smoothed distribution of the magnetic field, particles interact less with the mirrors if the initial $\mu$ is large ($\mu_{0}=0.8$).
@@
This scenario is quite different when scattering is more important (panels (a) and (b) of Figure 5). The initial pitch angle is less important for the overall behavior of the test particles and the identification of bounces with constant $\zeta$ is harder due to the presence of variations that hide the average behavior of the particle, introduced by pitch angle scattering. Comparing $\mu_{b}$ with its respective $\mu_{c,max}$ for this case shows a larger number of points below the orange line in Figure 5. Points identified here as bounce have already been compared with their neighbors, according to what was discussed in Section 3.1. These deviations can generally be explained by one of two reasons: either the particles are scattered and change directions or, due to scattering, $\mu$ later becomes smaller than $\mu_{c,max}$ along the particle trajectory and then the particle bounces while interacting with a mirror. In the former case, the misidentifying occurs due to our criterion not being strict enough. Still, the number of mirroring events are way more numerous than the events of bounces caused by scattering. In the latter case, we are not able to track the beginning of a bounce by looking at the peaks of $\mu$ versus time, so that defining a $\mu_{b}$ for these cases can be somewhat arbitrary. Still, these events are not frequent enough to interfere in our analysis.
@@
Figure 6 shows the PDFs of $\mu_{b}$ for all models. We calculate the PDF for all bounces integrated in time, after $t\times\Omega=2000$. Different colors represent different initial Larmor radii for the particles, with panels on the left-hand side showing models with an initial pitch angle such that $\mu_{0}=0.2$ and on the right-hand side with $\mu_{0}=0.8$. The top two panels show the results for models with the complete magnetic field whilst the bottom panels show results for models with the smoothed field as the background.
The distributions show that all models have a drop in probability for $\mu_{b}\gtrsim 0.6$. Also, most models show a peak in the distribution between $0.2$ and $0.3$.
@@
For the smoothed model r03mu80k5, only $\sim 50\%$ of the particles interact with a magnetic mirror, and about $1/5$ of the bouncing particles suffer only one or two bounces along the trajectory [^4] We note that in this case due to the lack of at least three bouncing events, we classified these events as mirroring. This introduces two points that need to be considered. First, the fraction of mirroring particles in this model is much smaller than in other models, where the pitch angle of the particle either is smaller at the beginning or resonant scattering changes the particle’s pitch angle enough that their probability of interacting with a mirror is higher. Second, as the particles travel inside the domain with higher values of $\mu$, they will see changes in the magnetic field and will only interact when the conditions for bounce are met, which will be skewed towards higher $\mu$ values due to the initial condition. The probability of a mirroring event occurring with $\mu>0.6$ is still small, but there are further complications that needs to be considered as we will discuss in the next section.
@@ 3.3 Particle diffusion in MHD turbulence
Figure 7: Top panels show the average squared displacement as a function of time normalized by the giro-frequency of the particles around the mean magnetic field. The second row shows the time derivative of the top panels, $\Delta\tilde{z}^{2}/\Delta t=\Delta<(z-z_{0})^{2}>/\Delta t$, giving the parallel diffusion coefficient when it becomes constant. The third row shows the average value of $\mu$ over all particles. The fourth and last row shows the average value of $\zeta/\zeta_{0}$ (Equation 3). Different colors represent different particles Larmor radii according to the legend. Time in x-axis is normalized by the initial value of the girofrequency $\Omega=c/R_{L}$ (eq. 2), with $c=1$ in code units.
@@
From top to bottom panels, Figure 7 compares the squared displacement of particles, the second panel shows the evolution of first derivative of the upper panel, $\Delta\tilde{z}^{2}/\Delta t$ (where we write $\tilde{z}^{2}=<(z-z_{0})^{2}>$ to concise the notation), the average value of $\mu$ over all particles, and the average value of $\zeta/\zeta_{0}$ (Equation 3) versus the normalized time $(t\times\Omega)$. The colors represent models with $R_{L}/L_{0}=0.03,0.06$ and $0.10$ for blue, orange and green, respectively. Note also that $\Omega$ is different for each Larmor radius. All models in this figure consider the complete magnetic field distribution of the background.
@@
The changes in $\Delta\tilde{z}^{2}/\Delta t$ over time in Figure 7 highlight the influence of the initial pitch angle. Initially, particles with low $\mu_{0}$ exhibit a smaller diffusion coefficient, with models starting with smaller Larmor radii requiring slightly more time to reach $<\mu>\simeq 0.5$. However, all the models incorporating the full magnetic field distribution demonstrate a stable diffusion process around $t\times\Omega\sim 10^{2}$. Note that when $\tilde{z}^{2}=<(z-z_{0})^{2}>$ increases at slower rate, the associated diffusion coefficient $D_{\parallel}=\Delta\tilde{z}^{2}/\Delta t$ becomes close to $\sim 10^{-2}$ for all models. However, there is actually a mild dependence of this saturation value with the Larmor radius of the model. This dependence is clearer in Figure 9 (bottom panel, Section 4), which does not use a logarithmic scale for the plot.
@@
Figure 8 depicts the same information as Figure 7, but for models featuring a smoothed distribution of the magnetic field as the background. The effects described earlier are more pronounced in these models. Specifically for $R_{L}/L_{0}=0.03$, there is a significant impact from magnetic mirrors on particles with $\mu_{0}=0.2$ as the influence of pitch-angle scattering diminishes. This is reflected in the time required for $\Delta\tilde{z}^{2}/\Delta t$ to stabilize. Towards the end of the observed period, $\Delta\tilde{z}^{2}/\Delta t$ is approximately one order of magnitude smaller for $\mu_{0}=0.2$ compared to $\mu_{0}=0.8$.
@@
The average behaviour of the adiabatic invariant in Figure 8 confirms that particles with smaller $\mu_{0}$ suffer mirroring for a longer duration initially. The average value of $<\zeta/\zeta_{0}>$ for models with $\mu_{0}=0.2$ begins to exhibit more significant variations only after $t\times\Omega=10^{2}$, whereas for models with $\mu_{0}=0.8$, variations commence after $t\times\Omega=10^{1}$.
@@
This behavior can be elucidated by considering the following. Particles do not interact with fluctuations smaller than their Larmor radius, and pitch angle scattering is associated with interactions with fluctuations having a wavelength $\lambda\sim R_{L}$. At the same time, large-scale perturbations induce the formation of magnetic mirrors. Referring back to Figure 1, $R_{L}/L_{0}=0.03$ corresponds to $k\sim 30$, which is indicative of magnetic energy smaller than $10^{-13}c.u.$. In essence, for $\mu_{0}=0.8$, particles can travel along magnetic field lines at much higher velocities, barely encountering mirrors initially due to their small pitch angle. However, as they traverse through small fluctuations, whether originating from numerical noise or residuals from smoothing, they experience scattering, initiating variations in $<\zeta/\zeta_{0}>$ earlier compared to the scenario when $\mu_{0}=0.2$. In the latter case, particles remain confined by a series of magnetic mirrors, predominantly diffusing perpendicular to the magnetic field.
@@
The effect of mirrors and scattering over the particles with smaller $R_{L}$ in Figures 7 and 8 is two-fold, as evidenced by the different magnetic field backgrounds. Mirrors can be very effective in confining particles in a system where scattering is suppressed at the same time that the particles can escape quickly in these same systems if they start their propagation in a direction close to parallel to the magnetic field. Scattering serves as an regulating mechanism, facilitating mirror diffusion along the magnetic field.
@@
An illustrative demonstration of this regulatory effect can be observed in the models depicted in Figure 8, with $R_{L}/L_{0}$ values of 0.06 and 0.10 (corresponding to $k\sim 17$ and $k\sim 10$, respectively). The temporal evolution of $\Delta\tilde{z}^{2}/\Delta t$ in these models closely resembles that of the scenario with a complete background magnetic field distribution (Figure 7), albeit with some nuanced distinctions. As the Larmor radius increases, the influence of the smoothing of the magnetic field distribution at small scales becomes less effective. Models with $R_{L}/L_{0}=0.06$ experience less scattering compared to those with $R_{L}/L_{0}=0.10$, yet the former ends up with a slightly higher diffusion coefficient than the latter. This seems to suggest that scattering becomes more efficient at higher energies, but we need to take into account that the dissipative scale for this background is around $\sim 0.2L_{0}$ (see Section 2.2). As the Larmor radius of the models becomes close to the dissipative scale, they can be scattered more efficiently even at higher $\mu$. This can also explain why the green models ($R_{L}/L_{0}=0.10$) seem to be better confined than the orange ones ($R_{L}/L_{0}=0.06$). At a small $\mu_{0}$, the non-resonant mirroring can well confine particles even when the Larmor radius is much smaller than the dissipative scale (see also Figure 10).
@@
Finally, Figures 7 and 8 also indicate that, although the particles can achieve a close to constant $\Delta\tilde{z}^{2}/\Delta t$ very quickly, which is indicative of a stable diffusion process, the entire process actually occurs over 5 to 10 times the size of the simulated box. This means that this initial phase of the diffusion, which can be heavily affected by mirrors, can occur over a considerable sized region around the source. Note however that this depends on how efficient resonant scattering is and it will also depend on the particle’s energy. Applications of these results and comparisons to observations will be discussed in the next section.

### v3

@@ 3.1 Particle trajectory and mirroring identification
In turbulence, particle trajectories are distorted by scattering. To distinguish such events from mirror diffusion, we introduce the following criterion. Let us recall that $\mu_{b}$ is the peak value of $\mu$ right before a bounce. We can thus evaluate the respective adiabatic invariant of $\mu_{b}$, $\zeta$ using Equation 3. This is shown in the third panels from top of Figures 3 and 4. For a particle with $n$ occurrences of $\mu=0$, and each peak occurring at times $t_{1}<t_{2}<...<t_{i}<...<t_{n}$, we compare $\zeta(t_{i})/\zeta_{avg}$ with $\zeta(t_{i-1})/\zeta_{avg}$ and $\zeta(t_{i+1})/\zeta_{avg}$ (to simplify notation, we consider $\zeta_{i}=\zeta(t_{i})/\zeta_{avg}$). With this, for a given bounce at time $t_{i}$, $\mu_{b}(t_{i})$ is discarded if $|\zeta_{i}-\zeta_{i-1}|>0.25$ and $|\zeta_{i}-\zeta_{i+1}|>0.25$. In other words, if the variation of $\zeta_{i}$ is bigger than $0.25$ compared to its neighbors, the bounce is discarded [^5] We note that the small oscillations in magnetic moment (as well as in $\mu$ and in $B$) are associated with gyrations, which are fundamentally different from the large stochastic variations caused by scattering. The threshold value $0.25$ was chosen to distinguish between these two effects.. For the first identified point, $\zeta_{1}$, we compare with the values of the next two identified points whilst for the last one, $\zeta_{n}$, we compare with the two before. If there are less than 3 bounce events, we simply accept the values of identified $\mu_{b}$.
For $R_{L}/L_{0}=0.03$ in the complete magnetic field background, over $73\%$ of bounces satisfy the condition defined above, $|(\zeta_{i}-\zeta_{i-1})/\zeta_{avg}|<0.25$. This proportion decreases slightly to $68\%$ and $67\%$, for $R_{L}/L_{0}=0.06$ and $0.10$, respectively. When the condition is tightened to $|(\zeta_{i}-\zeta_{i-1})/\zeta_{avg}|<0.2$, the corresponding percentages drop to $66\%$, $60\%$ and $59\%$ for $R_{L}/L_{0}=0.03$, $0.06$ and $0.10$, respectively. These results indicate the important role of magnetic mirroring in reducing the diffusion of particles. This point will be further explored in the following sections.
We have also tested an alternative method to identify bounces, which is described in Appendix B. We find that the results align with those described in this and the next section.
@@ 3.2 Pitch angle distribution of $\mu_{b}$
Figure 5: Comparison between the expected values of $\mu_{c,max}$ (Equation 6) and the identified bounce values from the simulations $\mu_{b}$. The orange line indicates where $\mu_{c,max}=\mu_{b}$. The color maps show the 2D probability density function normalized such that the sum over a bin value times the bin area is equal to 1. Only particles with more than two mirror interactions have been included. Top panels show the results for models with the complete magnetic field distribution, while the bottom panels show the results for models with the smoothed distribution of magnetic fields. On the left hand side we show models with $\mu_{0}=0.2$, while on the right hand side we show models with $\mu_{0}=0.8$.
Figure 6: Probability density distributions of $\mu_{b}$ (see text for further details). Colors indicate different Larmor radii of the particles. The top panels show the PDFs for models with the complete background magnetic field, whilst the bottom panels show PDFs for models with the smoothed magnetic field as background. On the left hand side we show models with $\mu_{0}=0.2$, while on the right hand side we show models with $\mu_{0}=0.8$. The distributions are independent of each other. The variance in the complete background model stays between 0.03 and 0.04. For the smoothed case stays between 0.02 and 0.03, except for the r15mu20k5 which has a variance closer to 0.01.
@@
Figure 5 compares the values of $\mu_{b}$ with the respective $\mu_{c,max}$ for the bounces identified in the models with $R_{L}=0.03L_{0}$. The color maps show the 2D probability density function normalized such that the sum over a bin value times the bin area is equal to 1. The top panels show the results for models with complete distribution of magnetic fields, while the bottom panels show the same for models that consider the smoothed distribution of magnetic fields. When $\mu_{0}$ is small, the number of bounces is generally higher than when $\mu_{0}$ is closer to 1. When the upper and bottom panels are compared, it is clear that the lack of scattering plays an important role in slowing down the diffusion parallel to the magnetic field, as evidenced by the models with a smoothed background.
@@
In the smoothed models in Figure 5 (panels (c) and (d)), we see that the identified $\mu_{b}$ events are mostly smaller than the corresponding predicted maximum values $\mu_{c,max}$. About $80-85\%$ of all the identified mirror events are above the orange line for all panels in Figure 5. This means that the particles are bouncing from larger-amplitude mirrors in the MHD cascade. As shown in Figure 4, the peaks maintain similar values while the particle bounces through successive mirrors. In addition, $\zeta$ stays almost constant during this process, indicating that $\mu_{b}$ is a good proxy for the bounce process. Note that for panel (d), particles with less than 3 mirror interactions are not included.
@@
This scenario is quite different when scattering is more important (panels (a) and (b) of Figure 5). The initial pitch angle is less important for the overall behavior of the test particles and the identification of bounces with constant $\zeta$ is harder due to the presence of variations that hide the average behavior of the particle, introduced by pitch angle scattering. Comparing $\mu_{b}$ with its respective $\mu_{c,max}$ for this case shows a larger number of points below the orange line in Figure 5. Points identified here as bounce have already been compared with their neighbors, according to what was discussed in Section 3.1. These deviations can generally be explained by one of two reasons: particles are scattered and change direction, or due to scattering, $\mu$ later becomes smaller than $\mu_{c,max}$ along the particle trajectory, and then the particle bounces while interacting with a mirror. In the former case, misidentifying occurs because our criterion is not strict enough. Still, the number of mirroring events is much more numerous than the events of bounces caused by scattering. In the latter case, we are not able to track the beginning of a bounce by looking at the peaks of $\mu$ versus time, so that defining a $\mu_{b}$ for these cases can be somewhat arbitrary. Still, these events are not frequent enough to interfere in our analysis [^7] We stress that mirror diffusion is present in both cases. If resonant scattering is weak, mirror diffusion will still act over the particles and can reduce the diffusion parallel to the magnetic field. In the smoothed case we also remove mirrors associated to small scales, which allows particles to travel further between mirrors..
The statistical significance of the region above the orange (critical) line in the plots of Figure 5 is clear. This line serves as a guide. As mentioned earlier, the identification of mirror interactions and scatterings in the turbulent magnetic fields is complex and makes it difficult to define where one effect ends and another begins. Nonetheless, the fact that most bounces occur with a nearly constant adiabatic invariant (Figures 3 and 4) and follow the maximum critical value for the interaction suggests that mirror diffusion is a significant effect and aligns with LX21 predictions.
@@
Figure 6 shows the PDFs of $\mu_{b}$ for all models. We calculate the PDF for all bounces integrated in time, after $t\times\Omega=4000$. Different colors represent different initial Larmor radii for the particles, with panels on the left-hand side showing models with an initial pitch angle such that $\mu_{0}=0.2$ and on the right-hand side with $\mu_{0}=0.8$. The top two panels show the results for models with the complete magnetic field whilst the bottom panels show results for models with the smoothed field as the background.
In these PDF distributions all models exhibit a drop in probability for $\mu_{b}\gtrsim 0.6$, along with a prominent peak between $0.2$ and $0.3$ in most cases. However, there is a noticeable scarcity of bounces identified with $\mu_{b}$ close to zero—not because such events are absent (see Figure 3), but because they are filtered out in the complete background models, where they appear as rapid fluctuations in $\mu$. In the smoothed background models, these near-zero $\mu_{b}$ bounces occur far less frequently. There is a correlation between $\mu$ and mirror size, as shown in Figure 6 in Zhang & Xu (2023). Small $\mu$ particles mainly interact with smaller mirrors, as they are more efficient in mirroring with larger longitudinal magnetic field gradient. In the smoothed case, where small mirrors are absent, the mirroring of small $\mu$ particles by larger mirrors is less efficient than in the complete case.
@@
For the smoothed model r03mu80k5, only $\sim 50\%$ of the particles interact with a magnetic mirror, and about $1/5$ of the bouncing particles suffer only one or two bounces along the trajectory [^8] We note that in this case, due to the absence of at least three bouncing events, these instances have been discarded from the analysis.. This introduces two points that need to be considered. First, the fraction of mirroring particles in this model is much smaller than in other models, where the pitch angle of the particle either is smaller at the beginning or resonant scattering changes the particle’s pitch angle enough that their probability of interacting with a mirror is higher. Second, as the particles travel inside the domain with higher values of $\mu$, they will see changes in the magnetic field and will only interact when the conditions for bounce are met, which will be skewed towards higher $\mu$ values due to the initial condition. The probability of a mirroring event occurring with $\mu>0.6$ is still small, but there are further complications that needs to be considered as we will discuss in the next section.
To complete the discussion in this section, in Table 2 we compare the total number of identified bounce events with the total number of candidates for all cases in Figure 6. Clearly, the percentage is larger for the smoothed models and for smaller particle Larmor radius. See also the discussion in Appendix B on how these percentages increase for an alternative method of identification of the bounce events. Nonetheless, the overall results remain consistent with the findings discussed above.
Table 2: Comparison between mirror events and the total number of candidates.
Complete $\mu_{0}=0.2$ $\mu_{0}=0.8$
$R_{L}$ Mirror Bounce Total % Mirror Bounce Total %
0.03 82636 123859 66.72 79733 120331 66,26
0.06 134473 223060 60.29 131082 218128 60.09
0.10 190035 322056 59.01 181022 307357 58.90
Smooth $\mu_{0}=0.2$ $\mu_{0}=0.8$
$R_{L}$ Mirror Bounce Total % Mirror Bounce Total %
0.03 265178 354073 74.89 114223 152117 75.09
0.06 113829 154984 73.45 102410 136214 75.18
0.10 90334 140372 64.35 87864 137993 63.67
@@ 3.3 Particle diffusion in MHD turbulence
Figure 7: Top panels show the average squared displacement.The second row shows the time derivative of the top panels, $\Delta\tilde{z}^{2}/\Delta t=\Delta<(z-z_{0})^{2}>/\Delta t$, giving the parallel diffusion coefficient when it becomes constant. The third row shows the average value of $\mu$ over all particles. The fourth and last row shows the average value of $\zeta/\zeta_{avg}$ (Equation 3). Different colors represent different particles Larmor radii according to the legend. Time in the x-axis is normalized by the initial value of the girofrequency $\Omega=c/R_{L}$ (eq. 2), with $c=1$ in code units.
@@
From top to bottom panels, Figure 7 compares the squared displacement of particles, the second panel shows the evolution of first derivative of the upper panel, $\Delta\tilde{z}^{2}/\Delta t$ (where we write $\tilde{z}^{2}=<(z-z_{0})^{2}>$ to concise the notation), the average value of $\mu$ over all particles, and the average value of $\zeta/\zeta_{avg}$ (Equation 3) versus the normalized time $(t\times\Omega)$. The colors represent models with $R_{L}/L_{0}=0.03,0.06$ and $0.10$ for blue, orange and green, respectively. Note also that $\Omega$ is different for each Larmor radius. All models in this figure consider the complete magnetic field distribution of the background.
@@
The evolution of $\Delta\tilde{z}^{2}/\Delta t$ over time in Figure 7 highlights the influence of the initial pitch angle. Initially, particles with low $\mu_{0}$ exhibit a smaller average squared displacement. The case with a smaller Larmor radius requires slightly more time to reach $<\mu>\simeq 0.5$. However, in all models that incorporate the full magnetic field distribution, the diffusion process stabilizes around $t\times\Omega\sim 10^{2}$. Notably, when $\tilde{z}^{2}=<(z-z_{0})^{2}>$ increases at a slower rate and increases linearly with time, the associated diffusion coefficient, $D_{\parallel}=\Delta\tilde{z}^{2}/\Delta t$, converges to approximately $10^{-2}$ across all models.
We note that the consistent asymptotic behavior across all models, with a finite diffusion mean free path, highlights the crucial role of mirroring. The primary function of resonant scattering in this process is to introduce stochastic variations in the pitch angle, rendering diffusion eventually independent of the initial pitch angle. At the same time, mirror diffusion ensures a finite mean free path, addressing the limitation of resonant scattering, which alone fails to reverse particle moving direction—commonly known as the 90-degree problem.
Additionally, there is a dependence of the value of $D_{\parallel}$ on the Larmor radius, which is more apparent in Figure 9 (bottom panel, Section 4), where a linear scale is used instead of a logarithmic one.
@@
Figure 8 depicts the same information as Figure 7, but for models featuring a smoothed distribution of the magnetic field as the background. The effects described earlier are more pronounced in these models. Specifically for $R_{L}/L_{0}=0.03$, there is a significant impact from magnetic mirrors on particles with $\mu_{0}=0.2$ as the influence of pitch-angle scattering diminishes. This is reflected in the time required for $\Delta\tilde{z}^{2}/\Delta t$ to stabilize. Towards the end of the observed period, $\Delta\tilde{z}^{2}/\Delta t$ is approximately one order of magnitude smaller for $\mu_{0}=0.2$ compared to $\mu_{0}=0.8$. Nevertheless, if we track the particles sufficiently long, this difference decreases to a factor two (see Figure 10 where the particles have evolved for a time 2.5 times longer). Therefore, one might speculate that after a much longer evolution, $\Delta\tilde{z}^{2}/\Delta t$ will eventually converge to the same value for both $\mu_{0}$, as in the complete case.
@@
The average behavior of the adiabatic invariant in Figure 8 confirms that particles with smaller $\mu_{0}$ suffer mirroring for a longer duration initially. The average value of $<\zeta/\zeta_{avg}>$ for models with $\mu_{0}=0.2$ begins to exhibit more significant variations only after $t\times\Omega=10^{2}$, whereas for models with $\mu_{0}=0.8$, variations commence after $t\times\Omega=10^{1}$.
@@
This behavior can be elucidated by considering the following. Particles marginally interact with fluctuations smaller than their Larmor radius, and pitch angle scattering is associated with interactions with fluctuations having a wavelength $\lambda\sim R_{L}$. At the same time, large-scale perturbations induce the formation of magnetic mirrors. In essence, for $\mu_{0}=0.8$, particles can travel along magnetic field lines at much higher velocities, barely encountering mirrors initially due to their small pitch angle. Essentially, the initial pitch angle is sufficiently small for the particles to undergo resonant scattering with the perturbations present in the system. The condition for this interaction will be discussed in more details in the next section. This process initiates variations in $<\zeta/\zeta_{avg}>$ earlier than in the scenario when $\mu_{0}=0.2$. In this latter case, particles remain confined by a series of magnetic mirrors, predominantly diffusing perpendicular to the mean magnetic field.
@@
The effect of mirrors and scattering over the particles with smaller $R_{L}$ in Figures 7 and 8 is two-fold, as evidenced by the different magnetic field backgrounds. Mirrors can effectively confine particles with large pitch-angles in systems where scattering is suppressed. However, particles may escape quickly if their initial propagation direction is nearly parallel to the magnetic field, as resonant scattering then becomes significant. At the same time, resonant scattering also acts as a regulating mechanism, enabling mirror diffusion as it causes changes in the pitch angle, allowing some particles to transit to a mirror diffusion dominated regime.
@@
An illustrative demonstration of this regulatory effect can be observed in the models depicted in Figure 8, with $R_{L}/L_{0}$ values of 0.06 and 0.10 (corresponding to $k\sim 17$ and $k\sim 10$, respectively). The temporal evolution of $\Delta\tilde{z}^{2}/\Delta t$ in these models closely resembles that of the scenario with a complete background magnetic field distribution (Figure 7), albeit with some nuanced distinctions. As the Larmor radius increases, the resonant condition is more readily satisfied, enabling more efficient pitch-angle scattering and causing the particle behavior to more closely resemble that observed in the complete background field case. Models with $R_{L}/L_{0}=0.06$ experience less scattering compared to those with $R_{L}/L_{0}=0.10$, yet the former ends up with a slightly higher diffusion coefficient than the latter. This seems to suggest that scattering becomes more efficient at higher energies, but we need to take into account that the dissipative scale for this background is around $\sim 0.2L_{0}$ (see Section 2.2). As the Larmor radius of the models becomes close to the dissipative scale, they can be scattered more efficiently even at higher $\mu$. This can also explain why the green models ($R_{L}/L_{0}=0.10$) seem to be better confined than the orange ones ($R_{L}/L_{0}=0.06$). At a small $\mu_{0}$, the non-resonant mirroring can well confine particles especially when the Larmor radius is much smaller than the dissipative scale. Though not very clear from Figure 8, it is important to note that the $R_{L}/L_{0}=0.03$ models (blue curves) also reach saturation at a later time. This is more evident in the bottom panels of Figure 10, where we have plotted a much larger time interval. Here, $<\Delta\tilde{z}^{2}>/\Delta t$ saturates for both $\mu_{0}=0.2$ and $0.8$, with the associated diffusion being at most one order of magnitude larger than the values reached in the complete models. Figure 11 illustrates the $\mu$ distribution near the end of the time interval presented in Figure 10. The results are shown for both smoothed models, with $\mu_{0}=0.2$ and $0.8$, and a Larmor radius of $R_{L}/L_{0}=0.03$. While the distribution is nearly constant and approaches isotropy, it does not become fully isotropic because of the inefficient scattering. The green curve (corresponding to $\mu_{0}=0.8$), for instance, shows a slightly excess of events close to $\mu=1.0$.
@@
Finally, Figures 7 and 8 also indicate that, although the particles can achieve a close to constant $\Delta\tilde{z}^{2}/\Delta t$ very quickly, which is indicative of a stable diffusion process, the entire process actually occurs over 5 to 10 times the size of the simulated box. This means that this initial phase of the diffusion, which can be heavily affected by mirrors, can occur over a considerably sized region around the source. Note however that this depends on how efficient resonant scattering is and it will also depend on the particle’s energy. Applications of these results and comparisons to observations will be discussed in the next section.